# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 5.2 Symmetries of lattice systems
Remark 5.1.
More generally, one could say that an almost local action of $G$ on a lattice system is a smooth homomorphism from $G$ to the group of Locally-Generated Automorphisms of this system. It was shown in [11] that LGAs form a diffeological group, thus the notion of a smooth homomorphism from a Lie group to the group of LGAs is well-defined. In applications, group actions on lattice systems are usually local in the more restrictive sense defined above, and to avoid a discussion of diffeological structures we adopt this more restrictive definition.
A local action of $G$ on $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ gives rise to a homomorphism from to $G$ the automorphism group of of ${\mathscr{A}}$ via $g\mapsto\otimes_{j\in\Lambda}{\rm Ad}_{\rho_{j}(g)}$ which is smooth on ${\mathscr{A}}_{a\ell}$. The corresponding generator ${\mathsf{Q}}$ is a homomorphism from ${\mathfrak{g}}$ to the Lie algebra of derivations of ${\mathscr{A}}_{a\ell}$ defined by
${\mathsf{Q}}:(a,{\mathcal{A}})\mapsto\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(a),{\mathcal{A}}],\quad a\in{\mathfrak{g}},\ {\mathcal{A}}\in{\mathscr{A}}_{a\ell},$ (98)
@@
${\mathsf{Q}}(f)({\mathcal{A}})=\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(f(j)),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}_{a\ell}.$ (99)
@@ 5.3 Equivariantization
As a preliminary step, for any $G$-invariant gapped state $\psi$ we are going to define a graded local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ which is a $G$-equivariant version of the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$. Recall that a graded local Lie algebra is a cosheaf of graded vector spaces and a pre-cosheaf of graded Lie algebras that satisfies the graded analogue of Property I. For example, if ${\mathfrak{F}}$ is a local Lie algebra and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally finite supercommutative graded algebra with finite-dimensional graded components $A_{k}$, [^12] A graded vector space is locally finite iff its graded components are finite-dimensional. then $U\mapsto{\mathfrak{F}}(U)\otimes A$ is a graded local Lie algebra. We denote it ${\mathfrak{F}}\otimes A$.
Fix a compact Lie group $G$ and distributive pre-ordered set $X$ and consider the category of graded local Lie algebras over $X$ equipped with a $G$-action. An object of this category is a graded local Lie algebra ${\mathfrak{F}}$ on which $G$ acts by automorphisms; morphisms are defined in an obvious manner. The first step is to define a functor ${\mathfrak{F}}\mapsto{\mathfrak{F}}^{G}$ from this category to the category of graded local Lie algebras over $X$ such that ${\mathfrak{F}}^{G}(U)$ is the Lie algebra of $G$-invariant elements of ${\mathfrak{F}}(U)$. It is clear how to define such a functor for coflasque pre-cosheaves of Lie algebras with Property I, but the pre-cosheaf ${\mathfrak{F}}^{G}$ will not be a cosheaf of vector spaces without further assumptions about ${\mathfrak{F}}$ and the $G$-action.
@@
$(g\cdot f)(j)={\rm Ad}_{g}f(j),\quad g\in G,j\in\Lambda.$ (101)
@@
$U\mapsto\left({\mathfrak{F}}(U)\otimes\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])\right)^{G}.$ (102)
@@ 6 Topological invariants of $G$-invariant gapped states
6.1 The commutator class of a pronilpotent acyclic pointed DGLA
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA, i.e. an acyclic DGLA with a distinguished central cycle ${\mathsf{B}}\in{\mathfrak{M}}_{-2}$. Assume it is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$.
@@ 6.1 The commutator class of a pronilpotent acyclic pointed DGLA
The commutator DGLA, denoted by $[{\mathfrak{M}},{\mathfrak{M}}]$, is defined to be the closure of the commutator subalgebra. Namely, ${\mathsf{q}}\in[{\mathfrak{M}},{\mathfrak{M}}]$ if and only if its projection is a linear combination of commutators in ${\mathfrak{M}}_{N}$ for all $N$.
Unlike ${\mathfrak{M}}$, the DGLA $[{\mathfrak{M}},{\mathfrak{M}}]$ is not necessarily acyclic. We would like to construct an obstruction to finding an element ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies $\partial{\mathsf{p}}={\mathsf{B}}$ and $[{\mathsf{p}},{\mathsf{p}}]=0$.
@@
$\partial{\mathsf{p}}+\frac{1}{2}[{\mathsf{p}},{\mathsf{p}}]={\mathsf{B}}.$ (103)
@@
Let ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then $[{\mathsf{p}},{\mathsf{p}}]$ is a cycle of the DGLA $[{\mathfrak{M}},{\mathfrak{M}}]$.
@@
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA which is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. Assume further that the structure morphisms $r_{N,N-1}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{N-1}$ are surjective and ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$ is central in ${\mathfrak{M}}_{N}$. Finally, assume $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$. Then ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. Furthermore, the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in $[{\mathfrak{M}},{\mathfrak{M}}]$ for ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is independent of the choice of ${\mathsf{p}}$.
A proof of this result can be found in Appendix B. It uses some results from deformation theory. We will call the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ the commutator class of the acyclic pointed DGLA $({\mathfrak{M}},{\mathsf{B}})$, for lack of a better name, and denote it ${\mathbf{com}}({\mathfrak{M}},{\mathsf{B}})$. The assignment $({\mathfrak{M}},{\mathsf{B}})\mapsto(H_{\bullet}([{\mathfrak{M}},{\mathfrak{M}}]),{\mathbf{com}}({\mathfrak{M}},{\mathsf{B}}))$ is a functor from the full sub-category of the category of acyclic pointed DGLAs to the category of pointed graded vector spaces ${{\mathsf{pVect}}_{\mathbb{Z}}}$.
@@ 6.2 Construction of topological invariants
The commutator class defined in the previous section is not a useful invariant of a gapped $G$-invariant state $\psi$ because it takes values in a set which itself depends on $\psi$. It is also not invariant under Locally Generated Automorphisms and thus is not an invariant of a gapped phase (see Remark 3.2). In this section we define a pairing between the commutator classes of ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and spherical CS cohomology classes of $S^{n-1}$ (the sphere at infinity) which takes values in the algebra of $G$-invariant symmetric polynomials on ${\mathfrak{g}}$. This gives a useful invariant of a gapped phase which is also an obstruction to promoting the global symmetry $G$ of the state to a local symmetry. Additionally, we will show that the invariant does not depends on the choice of the cover ${\mathfrak{U}}$ and thus is essentially unique.
Keeping in view further generalizations, we work over a sub-site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ which depends on an arbitray $W\in{\mathcal{C}\mathcal{S}}_{n}$. Let ${\hat{W}}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ be the image of $W$ under the equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. We continue to denote by ${\mathfrak{D}}_{al}$ the local Fréchet-Lie algebra which maps $U\in{{{\mathcal{C}\mathcal{S}}_{n}}/W}$ to ${\mathfrak{D}}_{al}$. For any cover ${\mathfrak{U}}$ of $W$ we denote by $\hat{\mathfrak{U}}$ the corresponding cover of $\hat{W}$.
@@
$(\partial^{*}\beta)_{i_{0}\dots i_{p}}=\sum_{k=0}^{p}(-1)^{k}\beta_{i_{0}\dots\widehat{i_{k}}\dots i_{p}}.$ (105)
@@
$\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$ (106)
@@
The above lemma remains true if one replaces ${\mathfrak{D}}_{al}$ with ${\mathfrak{D}}_{al}\otimes A$ where $A$ is any locally-finite graded vector space. Then the pairing $\langle{\mathsf{f}},\beta\rangle$ takes values in the space of inner derivations ${{\mathfrak{d}}_{al}}$ tensored with $A$.
@@
$\psi([{\mathsf{F}},{\mathsf{G}}])=\sum_{X\cap Y\neq\varnothing}\psi([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])=\sum_{Y}\psi({\mathsf{F}}({\mathsf{G}}^{Y}))=0,$ (107)
where the last equality follows from ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U)$. ∎
@@
$\left[C^{aug}_{\bullet+1}\left({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A\right),C^{aug}_{\bullet+1}\left({\mathfrak{U}},W,{\mathfrak{D}}^{\psi}_{al}\otimes A\right)\right].$ (108)
@@
$\langle{\mathsf{f}},\check{\delta}b\rangle=\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s}=-\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s},$ (109)
@@
The evaluations of cycles of the commutator DGLA on spherical CS cocycles can be used construct topological invariants of $G$-invariant gapped states on ${\mathbb{R}}^{n}$. We set $W={\mathbb{R}}^{n}$. The cycle we use as an input is the commutator class of $C_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ defined using the inhomogeneous Maurer-Cartan equation (Section 6.1). Since ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a sub-algebra of ${\mathfrak{D}}^{\psi}_{al}\otimes A$ with $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$, Prop. 6.3 applies to such cycles.
@@
By Prop. 4.7, the cohomology group $H^{l}_{CS}(S^{n-1},{\mathbb{R}})$ is one-dimensioal for $l=n-1$ and vanishes otherwise. It is natural to take the class of $({\mathfrak{U}},\beta)$ to be a generator of $H^{n-1}_{CS}(S^{n-1},{\mathbb{Z}})$. This generator is uniquely defined once an orientation of ${\mathbb{R}}^{n-1}$ has been chosen. Thus for any even $n$ we obtain an invariant of gapped $G$-invariant states on ${\mathbb{R}}^{n}$ taking values in $G$-invariant polynomials on ${\mathfrak{g}}$ of degree $(n+2)/2$, and this invariant changes sign when the orientation of ${\mathbb{R}}^{n}$ is changed. This is in an agreement with Chern-Simons field theory.

### v2

@@ 5.2 Symmetries of lattice systems
A local action of $G$ on $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ gives rise to a homomorphism from $G$ to the automorphism group of ${\mathscr{A}}$ via $g\mapsto\otimes_{j\in\Lambda}{\rm Ad}_{\rho_{j}(g)}$ which is smooth on ${\mathscr{A}}_{a\ell}$. The corresponding generator ${\mathsf{Q}}$ is a homomorphism from ${\mathfrak{g}}$ to the Lie algebra of derivations of ${\mathscr{A}}_{a\ell}$ defined by
${\mathsf{Q}}:(a,{\mathcal{A}})\mapsto\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(a),{\mathcal{A}}],\quad a\in{\mathfrak{g}},\ {\mathcal{A}}\in{\mathscr{A}}_{a\ell},$
@@
${\mathsf{Q}}(f)({\mathcal{A}})=\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(f(j)),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}_{a\ell}.$
@@ 5.3 Equivariantization
As a preliminary step, for any $G$-invariant gapped state $\psi$ we are going to define a graded local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ which is a $G$-equivariant version of the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$. Recall that a graded local Lie algebra is a cosheaf of graded vector spaces that is a pre-cosheaf of graded Lie algebras satisfying the graded analogue of Property I. For example, if ${\mathfrak{F}}$ is a local Lie algebra and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally finite supercommutative graded algebra with finite-dimensional graded factors $A_{k}$, [^10] A graded vector space is locally finite iff its graded components are finite-dimensional. then $U\mapsto{\mathfrak{F}}(U)\otimes A$ is a graded local Lie algebra. We denote it ${\mathfrak{F}}\otimes A$.
Fix a compact Lie group $G$ and a distributive pre-ordered set $X$ and consider the category of graded local Lie algebras over $X$ equipped with a $G$-action. An object of this category is a graded local Lie algebra ${\mathfrak{F}}$ on which $G$ acts by automorphisms; morphisms are defined in an obvious manner. The first step is to define a functor ${\mathfrak{F}}\mapsto{\mathfrak{F}}^{G}$ from this category to the category of graded local Lie algebras over $X$ such that ${\mathfrak{F}}^{G}(U)$ is the Lie algebra of $G$-invariant elements of ${\mathfrak{F}}(U)$. It is clear how to define such a functor for coflasque pre-cosheaves of Lie algebras with Property I, but the pre-cosheaf ${\mathfrak{F}}^{G}$ will not be a cosheaf of vector spaces without further assumptions about ${\mathfrak{F}}$ and the $G$-action.
@@
$(g\cdot f)(j)={\rm Ad}_{g}f(j),\quad g\in G,j\in\Lambda.$
@@
$U\mapsto\left({\mathfrak{F}}(U)\otimes\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])\right)^{G}.$
@@ 6 Topological invariants of $G$-invariant gapped states
6.1 The commutator class
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed DGLA, i.e. a DGLA with a distinguished central cycle ${\mathsf{B}}\in{\mathfrak{M}}_{-2}$. Assume it is a limit of an inverse system of nilpotent pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$.
@@ 6.1 The commutator class
The commutator DGLA, denoted by ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$, is defined to be the closure of the commutator subalgebra of ${\mathfrak{M}}$. Namely, ${\mathsf{q}}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ if and only if for any $N$ its projection to ${\mathfrak{M}}_{N}$ is a finite linear combination of commutators in ${\mathfrak{M}}_{N}$.
Even if ${\mathfrak{M}}$ is acyclic, the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ is not necessarily acyclic. We would like to construct an obstruction to finding an element ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies $\partial{\mathsf{p}}={\mathsf{B}}$ and $[{\mathsf{p}},{\mathsf{p}}]=0$.
@@
$\partial{\mathsf{p}}+\frac{1}{2}[{\mathsf{p}},{\mathsf{p}}]={\mathsf{B}}.$
@@
Let $({\mathfrak{M}},{\mathsf{B}})$ be pronilpotent pointed DGLA and ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then $[{\mathsf{p}},{\mathsf{p}}]$ is a cycle of the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$.
@@
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA which is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. Assume further that the structure morphisms $r_{N,N-1}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{N-1}$ are surjective and ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$ is central in ${\mathfrak{M}}_{N}$. Finally, assume $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$. Then ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. Furthermore, the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ for ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is independent of the choice of ${\mathsf{p}}$.
A proof of this result can be found in Appendix B. It uses some results from deformation theory. We will call the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ the commutator class of the acyclic pointed DGLA $({\mathfrak{M}},{\mathsf{B}})$, for lack of a better name, and denote it ${\mathbf{com}}({\mathfrak{M}},{\mathsf{B}})$. The assignment $({\mathfrak{M}},{\mathsf{B}})\mapsto(H_{\bullet}({\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}),{\mathbf{com}}({\mathfrak{M}},{\mathsf{B}}))$ is a functor from the full sub-category of the category of acyclic pointed DGLAs satisfying the conditions of Proposition 6.1, to the category ${{\mathsf{pVect}}_{\mathbb{Z}}}$ of pointed graded vector spaces.
@@ 6.2 Construction of topological invariants
The commutator class defined in the previous section is not a useful invariant of a gapped $G$-invariant state $\psi$ because it takes values in a set which itself depends on $\psi$. It is also not invariant under LGAs (defined in Section 3.2) and thus is not an invariant of a gapped phase (see Remark 3.2). In this section we define a pairing between the commutator classes of ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and spherical CS cohomology classes of the sphere at infinity $S^{n-1}$, which takes values in the algebra of $G$-invariant symmetric polynomials on ${\mathfrak{g}}$. This gives a useful invariant of a gapped phase which is also an obstruction to promoting the global symmetry $G$ of the state to a local symmetry. Additionally, we will show that the invariant does not depends on the choice of the cover ${\mathfrak{U}}$ and thus is essentially unique.
Keeping in view further generalizations, we work over a sub-site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ [^11] This is the so-called overcategory whose objects are equipped with a morphism to $W$. Due to coflasqueness, this amounts to a restriction to objects fuzzily included in $W$. which depends on an arbitray $W\in{\mathcal{C}\mathcal{S}}_{n}$. Let ${\hat{W}}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ be the image of $W$ under the equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. We continue to denote by ${\mathfrak{D}}_{al}$ the local Fréchet-Lie algebra which maps $U\in{{{\mathcal{C}\mathcal{S}}_{n}}/W}$ to ${\mathfrak{D}}_{al}$. For any cover ${\mathfrak{U}}$ of $W$ we denote by $\hat{\mathfrak{U}}$ the corresponding cover of $\hat{W}$.
@@
$(\partial^{*}\beta)_{i_{0}\dots i_{p}}=\sum_{k=0}^{p}(-1)^{k}\beta_{i_{0}\dots\widehat{i_{k}}\dots i_{p}}.$
@@
$\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$
@@
The above lemma remains true if one replaces ${\mathfrak{D}}_{al}$ with ${\mathfrak{D}}_{al}\otimes A$ where $A$ is any locally-finite graded vector space. Then the pairing $\langle{\mathsf{f}},\beta\rangle$ takes values in the space of inner derivations ${{\mathfrak{d}}_{al}}$ (Definition 3.3) tensored with $A$.
@@
$\psi([{\mathsf{F}},{\mathsf{G}}])=\sum_{X\cap Y\neq\varnothing}\psi([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])=\sum_{Y}\psi({\mathsf{F}}({\mathsf{G}}^{Y}))=0,$
where the last equality is because ${\mathsf{F}}$ preserves $\psi$. ∎
@@
$\overline{\left[C^{aug}_{\bullet+1}\left({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A\right),C^{aug}_{\bullet+1}\left({\mathfrak{U}},W,{\mathfrak{D}}^{\psi}_{al}\otimes A\right)\right]}.$ (24)
@@
$\langle{\mathsf{f}},\check{\delta}b\rangle=\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s}=-\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s},$
@@
The evaluations of cycles of the commutator DGLA on spherical CS cocycles can be used to construct topological invariants of $G$-invariant gapped states on ${\mathbb{R}}^{n}$. We set $W={\mathbb{R}}^{n}$. The cycle we use as an input is the commutator class of $C_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ defined using the inhomogeneous Maurer-Cartan equation (Section 6.1). Since ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a sub-algebra of ${\mathfrak{D}}^{\psi}_{al}\otimes A$ with $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$, Prop. 6.3 applies to such cycles.
@@
By Proposition 4.7, the cohomology group $H^{l}_{CS}(S^{n-1},{\mathbb{R}})$ is one-dimensioal for $l=n-1$ and vanishes otherwise. It is natural to take the class of $({\mathfrak{U}},\beta)$ to be a generator of $H^{n-1}_{CS}(S^{n-1},{\mathbb{Z}})$. This generator is uniquely defined once an orientation of ${\mathbb{R}}^{n-1}$ has been chosen. Thus for any even $n$ we obtain an invariant of gapped $G$-invariant states on ${\mathbb{R}}^{n}$ taking values in $G$-invariant polynomials on ${\mathfrak{g}}$ of degree $(n+2)/2$, and this invariant changes sign when the orientation of ${\mathbb{R}}^{n}$ is changed. This is in agreement with Chern-Simons field theory.